# Mismatch Quest: Visual and Textual Feedback for Image-Text Misalignment

### v1

@@ 7.1 Methodological Ablations Studies
Table 4: Comparing PaLI 5B models across configurations: baseline (untrained), fine-tuned for entailment, feedback, multitask with unfiltered data, and multitask (entailment+feedback). $+GD$ remarks that visual misalignment bounding-box prediction was performed in a 2-step way using the GroundingDino model. The study underscores the benefits of multitask training and the effectiveness of dataset filtering in enhancing performance.
We conduct an ablation study to evaluate the effectiveness of our methodologies. Initially, we compare our multi-task training approach, demonstrating superior performance with 75% entailment accuracy and a 0.72 BART-NLI [35] score in feedback, emphasizing the efficiency of our multi-task approach. Additionally, fine-tuning on our filtered dataset (77% of total data) improves performance in feedback and entailment tasks but degrades performance in others, highlighting the positive impact of our NLI model-based filtering. In a 2-step experiment simulating baselines, using GroundingDino [44] for grounding with predicted visual misalignment text labels improves bounding-box precision by 0.11 in F1-Score, showcasing its efficacy over our model.
@@ 7.2 Limitations and Future Work
  • No Visual Feedback: cases where no visual feedback is expected, as shown in Figure 7a. Our model is trained to always predict visual feedback and incorrectly generates it. To address this, we plan to enrich TV-Feedback with such scenarios, such as “an image of a horse” becoming “an image of a horse and a dog”, with feedback like “there is only a horse, not a dog and a horse”, and importantly, without generating a bounding box.
  • Multiple Misalignments: instances requiring identification of multiple misalignments, as shown in Figure 7b. Our model often detects only one issue where several exist. For enhancement, we will integrate such cases into TV-Feedback, like transforming “a white dog and a black cat” into “a white cat and a black dog”, with feedback addressing both the color and species misalignments and bounding boxes highlighting each.
  • Loose Bounding Boxes: As observed in Fig. 2 for the SD2.1 example, our model occasionally generates loose bounding boxes. For instance, rather than confining the bounding box to the pizza, it may encompass the entire person.
@@
Figure 7: Model limitations: (a) Misalignment due to a missing object, where no bounding box is expected, yet the model produces one over the only present object, a horse; (b) A scenario with multiple misalignments where the model identifies only one - here, the top parrot should be green (not white), and the bottom should be a white cat (not a green parrot). This limitation is evident when the model addresses only one error at a time, necessitating re-generation for complete correction.
@@ 8 Conclusion
Our work introduces an end-to-end approach that provides visual and textual feedback in text-to-image models, identifying and elucidating alignment discrepancies with visual annotations for targeted model refinement. We created TV-Feedback, a dedicated dataset for fine-tuning feedback models, leading to the development and evaluation of several robust models. We thoroughly tested on SeeTRUE-Feedback and out-of-distribution scenarios. While our current work focuses on improving feedback mechanisms within text-to-image models, we hope it will significantly contribute to the refinement of the generative accuracy of these models.
Acknowledgments
We express our gratitude to Amir Globerson, Jay Tenenbaum, and Shlomi Fruchter for their valuable feedback on earlier versions of the manuscript. We thank Jialin Wu and Sebastian Goodman for their help. This research was supported in part by the Israel Science Foundation (grants no. 2492/20, 3441/21 and 3611/21), Len Blavatnik and the Blavatnik family foundation.
@@ Acknowledgments
Appendix A Additional Results
@@
Figure 8: Additional qualitative comparison examples from SeeTRUE-Feedback. The PaLI 55B model, fine-tuned on TV-Feedback, produces the most aligned and concise feedback compared to the rest of the models.
@@
Appendix B ConGen-Feedback Details
As detailed in the paper, to generate our training data using ConGen-Feedback, we collected aligned text-image pairs from six different datasets: PickaPic [33], ImageReward [74], COCO [39], Flickr30k [53], Open Images [34], and ADE20k [84]. It’s important to note that we selected only the positive pairs from these datasets, totaling 1,302,201 text-image pairs. From these pairs, we generated negative examples with misalignments accompanied by textual and visual feedback. To validate the generation and maintain a high quality of generated data, we filtered out samples based on textual entailment tests, as detailed in Section B.2, resulting in 1,087,912 text-image pairs for our filtered ConGen-Feedback version.
B.1 LLM Prompts
@@ B.2 Generation Validation
Feedback Evaluation Task Template: Contradiction Correctness: Premise: $OriginalCaption$ Hypothesis: $GeneratedContradiction$ Feedback Correctness: Premise: EXPECTED CAPTION: $GeneratedContradiction$ ACTUAL CAPTION: $OriginalCaption$ Hypothesis: $GeneratedFeedback$
Example A: Contradiction Correctness: Original Caption: “A black cat playing on top of a wooden chair.” Generated Contradiction: “A black cat playing underneath a wooden chair” Generated Feedback: “The cat is on top of the chair, not underneath it.” Contradiction Correctness score: 0.02 ✓ Feedback Correctness score: 0.97 ✓
Example B: Contradiction Correctness: Original Caption: “The man is riding along the beach in a two-wheeled cart pulled by a horse.” Generated Contradiction: “The man is driving along the beach in a two-wheeled cart pulled by a horse.” Generated Feedback: “The man is riding along the beach, not driving” Contradiction Correctness score: 0.7 ✗ Feedback Correctness score: 0.95 ✓
Example C: Contradiction Correctness: Original Caption: “a dining room table with a bottle of wine and wine glasses and a pot of carrots” Generated Contradiction: “a dining room table with a bottle of wine and wine glasses and a pot of potatoes” Generated Feedback: “The carrots are replaced with potatoes.” Contradiction Correctness score: 0.06 ✓ Feedback Correctness score: 0.01 ✗
@@
Appendix C SeeTRUE-Feedback Details
@@
Similar to the approach detailed in B.1, we utilize PaLM 2 to generate the required labels. In this case, there is no need to generate a contradiction and feedback for a sample. Instead, we need to summarize the three human feedbacks into a single and concise feedback and generate the remaining labels, such as the misalignment cue source in the caption and the textual label for the visual bounding box.
@@
Appendix D Training details
@@
Hardware Requirements: The computational resources required for fine-tuning vary depending on the specific PaLI model. PaLI 5B and 17B fine-tuning require four v4 chips [32], while fine-tuning for PaLI 55B necessitates sixteen v4 chips.
@@
For the CAPTION description, ensure it thoroughly represents the true content of the CAPTION (e.g., ”CAPTION: a blue ball on a shelf”).
For the CONTRADICTION description, keep it brief and directly replicate the relevant segment (e.g., ”CONTRADICTION: green ball”).
@@
CAPTION: ”A person dressed as Captain America is holding a shield in front of a wall.”
@@
CAPTION: ”a close up of a lion with a long mane sitting on the grass with trees in the background”
@@
CAPTION: ”A painting of a man standing in front of a fire in a field with trees and sky in the background .”
@@
CAPTION: ”A tattooed woman in a purple brassiere and purple underpants is standing in front of a table with a bottle on it.”
@@
CAPTION: ”portrait of a smiling man and woman hugging each other on the street with trees and buildings in the background”
@@
For the CAPTION description, ensure it thoroughly represents the true content of the CAPTION (e.g., ”CAPTION: a blue ball on a shelf”).
For the CONTRADICTION description, keep it brief and directly replicate the relevant segment (e.g., ”CONTRADICTION: green ball”).
@@
CAPTION: ”a waterfall in a tropical forest with green plants and trees”
@@
CAPTION: ”a cartoon of a red haired mermaid holding a fish in the sea”
@@
SUMMARIZED FEEDBACK: Given a caption and a set of visual feedbacks, your objective is to create a concise and informative feedback statement that captures the most significant misalignment between the caption and the visual content. No need to mention ”at the caption” or ”at the image”.
@@
CAPTION: ”A woman with a blue shirt and yellow flower headband, sitting on a wooden bench outside on the grass.”
FEEDBACKS: [”The woman’s shirt is yellow, not blue.”, ”The woman sitting on the wooden bench is wearing a white shirt, not a blue shirt.”, ”The woman has a yellow shirt on, not blue.”]
@@
FEEDBACKS: [”A couch is to the left of a coffee table, there is no chair present.”, ”A couch and two tables under a window, there is no chair.”, NaN]
@@
FEEDBACKS: [”A sheep is facing the camera, not to the right of a wine glass, which is not present.”, ”A sheep looking at the camera, there is no wine glass.”, ”There is only a sheep alone in a field, it is not to the right of a wine glass”]
@@
FEEDBACKS: [NaN, ”A pumpkin has been put on top of a kitchen stool. There is a fork sticking into the pumpkin. The scene is viewed from above. There is no tomato on top of the pumpkin.”, ”There is only a pumpkin, and no tomato”]
@@
FEEDBACKS: [”The glass ball is blue, not red.”, ”The ball with the reflection should be blue instead of red.”, ”The ball is blue, not red”]
@@
FEEDBACKS: [”There is no meat present, nor are there any chopsticks, simply the bowl with fruit and vegetable.”, ”There is a bowl of food, but no chopsticks.”, ”There is a salad with fruit and vegetables in a yellow bowl but there does not appear to be meat in it and there are no chopsticks.”]
@@
FEEDBACKS: [”The wall is gray, not white.”, ”A bucket has two canes in it, not umbrellas and canes and it is front of a gray wall, not a white wall.”, ”A cup has canes in it, but no umbrellas”]
@@
FEEDBACKS:[”A small white toilet is here without any rolls of toilet paper, not any spares.”, ”There is a toilet but no toilet paper.”, ”The toilet only has curtains/walls next to it, not spare rolls of toilet paper.”]
@@
FEEDBACKS: [”A dog is holding a frisbee in its mouth”, ”A dog is holding a frisbee in its mouth.”, ”A dog is holding a frisbee, not a cat”]
@@
FEEDBACKS: [”A dog is holding a frisbee in its mouth”, ”A dog is holding a frisbee in its mouth.”, ”A dog is holding a frisbee, not a cat”]

### v2

@@ 7.1 Methodological Ablations Studies
We conduct an ablation study to evaluate our methodologies. Our multi-task training approach achieves superior performance, with 75% entailment accuracy and a 0.72 BART-NLI [35] score in feedback, highlighting its efficiency. Fine-tuning on our filtered dataset (77% of total data) improves feedback and entailment tasks but degrades others, underscoring the positive impact of NLI model-based filtering. In a 2-step experiment simulating baselines, using GroundingDino [44] for grounding with predicted visual misalignment text labels improves bounding-box precision by 0.11 in F1-Score, showcasing its efficacy over our model.
Table 4: Comparing PaLI-3 [12] models: baseline, fine-tuned (entailment, feedback), multitask (unfiltered data, entailment+feedback). $+GD$ denotes two-step visual misalignment (b-box) prediction via GroundingDino. The study underscores the benefits of multitask training and the effectiveness of dataset filtering in enhancing performance.
@@ 7.2 Limitations and Future Work
  • No Visual Feedback: In cases where no visual feedback is expected (Figure 8a), our model incorrectly predicts it. To address this, we plan to enrich TV-Feedback with scenarios like “an image of a horse” becoming “an image of a horse and a dog,” with feedback like “there is only a horse, not a dog and a horse,” and without generating a bounding box.
  • Multiple Misalignments: Instances requiring identification of multiple misalignments (Figure 8b). Our model often detects only one issue where several exist. We will enhance TV-Feedback with cases like transforming “a white dog and a black cat” into “a white cat and a black dog,” with feedback addressing both color and species misalignments and bounding boxes highlighting each. In Appendix 0.A , we show how a finetuned model detects multiple misalignments sequentially using a MagicBrush [83] dataset example. The model identifies one misalignment at a time, and feedback signals guide an instruction editing model to iteratively correct them.
  • Loose Bounding Boxes: As observed in Fig. 2 for the SD2.1 example, our model occasionally generates loose bounding boxes. For instance, rather than confining the b-box to the pizza, it may encompass the entire person.
@@
Figure 8: Model limitations: (a) Misalignment due to a missing object, where the model incorrectly adds a bounding box over a horse; (b) Multiple misalignments, with the model only identifying one - the top parrot should be green and the bottom a white cat. The model requires multiple iterations for full correction.
@@ 8 Conclusion
Our research develops an end-to-end strategy providing visual and textual feedback for text-to-image models, targeting and clarifying alignment issues for refinement. We introduced TV-Feedback, a specialized dataset for fine-tuning feedback in these models, leading to several robust developments. This dataset and methodology demonstrate broad potential, notably in enhancing text-to-image generation, improving dataset annotation accuracy, and refining image captioning through detailed feedback. Our comprehensive testing on SeeTRUE-Feedback and various scenarios validates our approach’s effectiveness. While primarily aimed at text-to-image feedback enhancement, we anticipate our work will significantly improve generative model accuracy across different domains.
@@
Appendix 0.A Additional Results
@@
As illustrated in Fig. 12, we show how a fine-tuned model can sequentially detect multiple misalignments using an example from the MagicBrush dataset. While the model identifies misalignments one at a time, the feedback signals direct an instruction editing model to iteratively correct these errors. This method can be used as a baseline for future research addressing multiple misalignments, particularly in the context of long image descriptions.
Figure 9: Additional qualitative comparison examples from SeeTRUE-Feedback. The PaLI-X model, fine-tuned on TV-Feedback, produces the most aligned and concise feedback compared to the rest of the models.
@@
Figure 12: Demonstrating the iterative correction setup of image-text misalignments using a fine-tuned model on an example from the MagicBrush test dataset. In each iteration, the predicted feedback signals guide the model to sequentially correct the description until it fully aligns with the image.
Appendix 0.B ConGen-Feedback Details
As detailed in the paper, to generate our training data using ConGen-Feedback, we collected aligned text-image pairs from six different datasets: PickaPic [33], ImageReward [74], COCO [39], Flickr30k [53], Open Images [34], and ADE20k [85]. It’s important to note that we selected only the positive pairs from these datasets, totaling 1,302,201 text-image pairs. From these pairs, we generated negative examples with misalignments accompanied by textual and visual feedback. To validate the generation and maintain a high quality of generated data, we filtered out samples based on textual entailment tests, as detailed in Section 0.B.2, resulting in 1,087,912 text-image pairs for our filtered ConGen-Feedback version.
0.B.1 LLM Prompts
@@ 0.B.2 Generation Validation
Feedback Evaluation Task Template:
Contradiction Correctness:
Premise: $OriginalCaption$
Hypothesis: $GeneratedContradiction$
Feedback Correctness:
Premise: EXPECTED CAPTION: $GeneratedContradiction$ ACTUAL CAPTION: $OriginalCaption$
Hypothesis: $GeneratedFeedback$
Example A:
Contradiction Correctness:
Original Caption: “A black cat playing on top of a wooden chair.”
Generated Contradiction: “A black cat playing underneath a wooden chair”
Generated Feedback: “The cat is on top of the chair, not underneath it.”
Contradiction Correctness score: 0.02 ✓
Feedback Correctness score: 0.97 ✓
Example B:
Contradiction Correctness:
Original Caption: “The man is riding along the beach in a two-wheeled cart pulled by a horse.”
Generated Contradiction: “The man is driving along the beach in a two-wheeled cart pulled by a horse.”
Generated Feedback: “The man is riding along the beach, not driving”
Contradiction Correctness score: 0.7 ✗
Feedback Correctness score: 0.95 ✓
Example C:
Contradiction Correctness:
Original Caption: “a dining room table with a bottle of wine and wine glasses and a pot of carrots”
Generated Contradiction: “a dining room table with a bottle of wine and wine glasses and a pot of potatoes”
Generated Feedback: “The carrots are replaced with potatoes.”
Contradiction Correctness score: 0.06 ✓
Feedback Correctness score: 0.01 ✗
@@
Appendix 0.C SeeTRUE-Feedback Details
@@
Similar to the approach detailed in 0.B.1, we utilize PaLM 2 to generate the required labels. In this case, there is no need to generate a contradiction and feedback for a sample. Instead, we need to summarize the three human feedbacks into a single and concise feedback and generate the remaining labels, such as the misalignment cue source in the caption and the textual label for the visual bounding box.
@@
Appendix 0.D Evaluation Metrics - Additional Details
This section builds upon Section 5.2 by providing deeper and more technical details of the evaluation process and metrics calculation reported in Table 2. We compare our models against current leading VLM baselines using several metrics, which are explained in detail in this section. It is important to emphasize that our fine-tuned models deliver all required results in an end-to-end manner, whereas for the baselines, we must generate the answer using a different prompt each time.
To evaluate the quality of textual outputs, we conducted human annotation study as mentioned in Section 6 and automatic metric calculation. Figure 15 we show the annotators agreement for the feedback predictions correctness produced by each model. For the automatic metric to evaluate we opt for BART NLI [35]. Based on Figure 7, we find it to be the most correlated metric with human ratings. The NLI algorithm receives two textual inputs: a ‘premise’ and a ‘hypothesis’, and outputs an entailment score ranging from 0 to 1, indicating semantic alignment (a higher score implies greater semantic alignment). In our experiments, we consider the ground truth text as the ‘premise’ and the model prediction as the ‘hypothesis’. Figure 5 in the main paper presents examples with metrics calculations results.
Feedback NLI (Textual Feedback Quality): To obtain feedback from the baseline models, we query them with the question: “What are the misalignments between the image and the text: <text>?” As mentioned earlier, we use BART NLI to measure the semantic alignment between the ground truth and the predicted feedback answer.
Textual Misalignment NLI: In this task, we aim to guide the VLM to identify which part of the input caption causes misalignment between the image and the text. We pose the question to the model as follows: “Which part of <text> does not align with the image” Again, we utilize the BART NLI model to assess the semantic quality of the predictions compared to the ground truth.
Visual Misalignment - F1-Score@0.75: This metric aims to assess the VLM’s accuracy in marking the visual source of misalignment between the image and the caption. The visual misalignment is represented by a labeled bounding box. Since baseline VLMs cannot predict a bounding box, we divide this process into two steps. Initially, we query the VLM with: “Which part of the following image does not align with the <text>?”. To obtain a bounding box corresponding to the predicted visual misalignment, we employ the GroundingDino [44] grounding model to extract a bounding box from the image, guided by the predicted label. Our fine-tuned models, trained with TV-Feedback data, do not require this two-step process and can predict feedback, including bounding boxes, in a single inference step. In Table 2, we report the performance of our fine-tuned models using the same pipeline as the VLM baseline, as well as results from the PaLI models. The chosen metric for comparing model performance is the F1-Score@0.75, which sets the IoU threshold at 0.75 and balances precision and recall to measure the quality of localization.
Binary Class Accuracy (Image-Text Alignment): This metric compares the models’ abilities to determine whether there is any misalignment between an image and its caption in a binary fashion. The input query for the VLMs is: “Does this image entail the description <text>?”, expecting ‘yes’ or ‘no’ answers. The reported metric is the average accuracy across the SeeTRUE test dataset.
Figure 15: Bar chart showing consensus levels among three annotators for model feedback predictions, segmented by the number (0-3) of raters in agreement. For example, in the bottom bar, all three raters agreed on prediction’s correctness in 67.4% of cases, while for LLaVA, this occurred in only 17.1% of cases.
Appendix 0.E Training details
@@
Hardware Requirements: The computational resources required for fine-tuning vary depending on the specific PaLI model. PaLI and PaLI-3 fine-tuning require four v4 chips [32], while fine-tuning for PaLI-X necessitates sixteen v4 chips.
@@
For the CAPTION description, ensure it thoroughly represents the true content of the CAPTION (e.g., "CAPTION: a blue ball on a shelf").
For the CONTRADICTION description, keep it brief and directly replicate the relevant segment (e.g., "CONTRADICTION: green ball").
@@
CAPTION: "A person dressed as Captain America is holding a shield in front of a wall."
@@
CAPTION: "a close up of a lion with a long mane sitting on the grass with trees in the background"
@@
CAPTION: "A painting of a man standing in front of a fire in a field with trees and sky in the background ."
@@
CAPTION: "A tattooed woman in a purple brassiere and purple underpants is standing in front of a table with a bottle on it."
@@
CAPTION: "portrait of a smiling man and woman hugging each other on the street with trees and buildings in the background"
@@
For the CAPTION description, ensure it thoroughly represents the true content of the CAPTION (e.g., "CAPTION: a blue ball on a shelf").
For the CONTRADICTION description, keep it brief and directly replicate the relevant segment (e.g., "CONTRADICTION: green ball").
@@
CAPTION: "a waterfall in a tropical forest with green plants and trees"
@@
CAPTION: "a cartoon of a red haired mermaid holding a fish in the sea"
@@
SUMMARIZED FEEDBACK: Given a caption and a set of visual feedbacks, your objective is to create a concise and informative feedback statement that captures the most significant misalignment between the caption and the visual content. No need to mention "at the caption" or "at the image".
@@
CAPTION: "A woman with a blue shirt and yellow flower headband, sitting on a wooden bench outside on the grass."
FEEDBACKS: ["The woman’s shirt is yellow, not blue.", "The woman sitting on the wooden bench is wearing a white shirt, not a blue shirt.", "The woman has a yellow shirt on, not blue."]
@@
FEEDBACKS: ["A couch is to the left of a coffee table, there is no chair present.", "A couch and two tables under a window, there is no chair.", NaN]
@@
FEEDBACKS: ["A sheep is facing the camera, not to the right of a wine glass, which is not present.", "A sheep looking at the camera, there is no wine glass.", "There is only a sheep alone in a field, it is not to the right of a wine glass"]
@@
FEEDBACKS: [NaN, "A pumpkin has been put on top of a kitchen stool. There is a fork sticking into the pumpkin. The scene is viewed from above. There is no tomato on top of the pumpkin.", "There is only a pumpkin, and no tomato"]
@@
FEEDBACKS: ["The glass ball is blue, not red.", "The ball with the reflection should be blue instead of red.", "The ball is blue, not red"]
@@
FEEDBACKS: ["There is no meat present, nor are there any chopsticks, simply the bowl with fruit and vegetable.", "There is a bowl of food, but no chopsticks.", "There is a salad with fruit and vegetables in a yellow bowl but there does not appear to be meat in it and there are no chopsticks."]
@@
FEEDBACKS: ["The wall is gray, not white.", "A bucket has two canes in it, not umbrellas and canes and it is front of a gray wall, not a white wall.", "A cup has canes in it, but no umbrellas"]
@@
FEEDBACKS:["A small white toilet is here without any rolls of toilet paper, not any spares.", "There is a toilet but no toilet paper.", "The toilet only has curtains/walls next to it, not spare rolls of toilet paper."]
@@
FEEDBACKS: ["A dog is holding a frisbee in its mouth", "A dog is holding a frisbee in its mouth.", "A dog is holding a frisbee, not a cat"]
@@
FEEDBACKS: ["A dog is holding a frisbee in its mouth", "A dog is holding a frisbee in its mouth.", "A dog is holding a frisbee, not a cat"]